# AdaAugment: A Tuning-Free and Adaptive Approach to Enhance Data Augmentation

### v2

@@ 4.3 Results on ImageNet
Table 3: Top-1 accuracy (%) on ImageNet dataset.
@@
baseline HaS Cutout CutMix GM Mixup AA FAA RA DADA TA AdaAugment
77.1 77.2 77.1 77.2 77.9 77.0 77.6 77.6 77.6 77.5 77.9 78.2
We evaluate AdaAugment on large-scale IamgeNet dataset (Krizhevsky et al., 2017).
4.4 Transfer Learning
@@
Transfer learning plays a pivotal role in assessing the models’ transferability (Zhuang et al., 2021), a common practice in data augmentation (Gong et al., 2021; Yang et al., 2023). To quantitatively assess the efficacy of various DA methods, we assess the transferred accuracy of models using different DA methods. Specifically, we pre-train ResNet-50 models on the CIFAR-100 dataset utilizing different augmentations, followed by a subsequent fine-tuning of these models on the CIFAR-10 dataset. Thus, superior DA techniques lead to higher transferred test accuracy.
@@
In Fig. 4, we show the transferred accuracy achieved by various DA methods. While the discrepancies in transferred accuracy may appear modest, it is noteworthy that AdaAugment consistently outperforms others. This consistent and effective performance underscores our method’s capability to enhance model transferability performance.
Figure 4: Transferred test accuracy of various DA methods.
4.5 Analytical Results
4.5.1 Visualization of Effectiveness
Since the objective of DA is to enhance the generalization performance of models, in this section, we conduct a comparative analysis of effectiveness visualization between models utilizing AdaAugment and those without. Specifically, we leverage the t-SNE algorithm to visualize the feature embedding of the MNIST test set, utilizing feature maps derived from different trained models. Theoretically, models that showcase more robust generalization capabilities tend to excel in feature extraction ability. Fig. 5 presents the t-SNE visualization results for the baseline model and the model trained with AdaAugment. Compared to Fig. 5(a), it can be seen that Fig. 5(b) exhibits an optimized geometric structure characterized by enhanced inter-cluster separation and intra-cluster compactness.
Furthermore, to furnish a quantitative evaluation of these embedding results, we employ the Dunn Index (DI) (Ncir et al., 2021) as an evaluative metric. The DI is mathematically defined as follows:
$DI=\frac{\min_{1\leq i\neq j\leq m}\delta\left(C_{i},C_{j}\right)}{\max_{1\leq j\leq m}\Delta_{j}}.$ (5)
Here, separation $\delta\left(C_{i},C_{j}\right)$ denotes the inter-cluster distance metric between clusters $C_{i}$ and $C_{j}$, and compactness $\Delta_{j}$ calculates the mean distance between all pairs within each cluster. A higher DI value signifies superior clustering. In Fig. 5, the DI values for the baseline and AdaAugment are $5.03\times 10^{-5}$ and $1.05\times 10^{-4}$, respectively. Consequently, AdaAugment achieves a DI value that is 108.7% higher than the baseline, demonstrating its efficacy in enhancing model performance. Based on both quantitative and qualitative analysis, we validate the substantial effectiveness of AdaAugment in improving model performance.
@@ 4.5.1 Visualization of Effectiveness
Figure 5: Visualization of effectiveness using t-SNE algorithm. DI: Dunn index.
4.5.2 Complexity Analysis
Table 4: The additional architecture parameters and training time of the auxiliary policy network. We report average $\pm$ std.
@@ 4.5.2 Complexity Analysis
Most existing DA methods typically rely on predefined random magnitude values during training. Although this mechanism incurs minimal additional computational overhead in online training, it may not be optimally aligned with the dynamic evolution of deep models, potentially introducing undesirable side effects. To address this, AdaAugment incorporates an auxiliary policy network into online training to determine the magnitudes of DA operations dynamically. Although this adjustment inevitably introduces slightly additional parameters and computational complexity, it is a deliberate trade-off to achieve more effective DA.
@@
In this section, we thoroughly analyze the parameter and time complexity of AdaAugment to assess its efficiency. Since the policy network’s parameter complexity relates to the target network’s feature space, we report the parameter complexity on two different deep networks, i.e., ResNet-18 and WRN-28-10. In Table 4, we present the complexity analysis for AdaAugment using the CIFAR-10 dataset, following the methodology employed in prior studies (Cubuk et al., 2019; Lim et al., 2019; Lin et al., 2023a). It can be observed that the policy network’s parameter complexity results in a mere 1.3% increase for ResNet-18 (with 11.7M parameters) and a 0.52% increase for WRN-28-10 (with 36.5M parameters). Furthermore, the overall training costs experience an increase of less than 0.5 GPU hours for these networks. These findings highlight the negligible training overhead of AdaAugment.
Notably, despite the marginal increases in training costs, AdaAugment delivers substantial accuracy improvements, achieving a notable enhancement of 1.47% and 2.04% for ResNet-18 and WRN-28-10, respectively. This improvement underscores AdaAugment’s remarkable effectiveness-efficiency trade-off, demonstrating its capacity to enhance model performance while incurring minimal training costs.
4.5.3 Convergence Analysis
In this section, we conduct a convergence analysis of AdaAugment throughout the entire training process and compare it against two of the most representative DA methods, AutoAugment and TrivialAugment. Specifically, to present the dynamic evolution of test errors during training, we train ResNet-18 models on CIFAR-10 using a multi-step learning rate decay schedule. The learning rate is initialized as 0.1 and multiplied by 0.2 at epochs 60, 120, 160, 220, and 280. From Fig. 6, it can be observed that AdaAugment significantly enhances model performance, particularly after the second learning rate drop. Moreover, even after the initial learning rate drop, AdaAugment consistently maintains lower error rates than other methods. These empirical findings not only underscore the superior efficacy of AdaAugment in terms of model performance enhancement but also emphasize its capacity to expedite the convergence of the model during training.
@@ 4.5.3 Convergence Analysis
Figure 6: Convergence analysis of training process.
5 Conclusion

### v3

@@ IV-B Results on Tiny-ImageNet
TABLE VI: Top-1 accuracy (%) on ImageNet-1k dataset with ResNet-50. Some results are adapted from [23, 28, 25].
@@
baseline HaS [45] FAA [46] DADA [28] Cutout [14] CutMix [18] MADAug [50] GridMask [15] AA [21] RA [22] TeachAug [25] TA [23] RE [43] AdaAugment
77.1 77.2 77.6 77.5 77.1 77.2 78.3 77.9 77.6 77.8 78.0 77.9 77.3 78.3
TABLE VII: Performance (%) on ImageNet-1k with ViT-Base, ViT-Large, and Swin-Transformer using a 4-V100-GPU server.
Model baseline AdaAugment
ViT-B 81.46 82.55
ViT-L 83.50 84.67
Swin-Transformer 85.00 85.65
TABLE VIII: Transferred test accuracy (%) on CIFAR-10 of various DA methods. The pretrained ResNet-50 model is trained on both CIFAR-100 (upper row) and Tiny-ImageNet (bottom row).
baseline HaS [45] FAA [46] DADA [28] Cutout [14] CutMix [18] MADAug [50] GridMask [15] AA [21] RA [22] TeachAug [25] TA [23] RE [43] AdaAugment
91.53$\pm$.03 92.51$\pm$.24 92.28$\pm$.13 92.58$\pm$.09 92.42$\pm$.20 92.81$\pm$.47 92.84$\pm$.10 91.49$\pm$.10 92.82$\pm$.04 92.78$\pm$.23 92.83$\pm$.18 92.80$\pm$.16 92.55$\pm$.05 93.06$\pm$.25
64.02$\pm$.05 66.84$\pm$.06 70.32$\pm$.63 69.04$\pm$.43 65.54$\pm$.75 69.29$\pm$.09 72.82$\pm$.32 64.88$\pm$.43 69.53$\pm$.53 64.68$\pm$.97 69.98$\pm$.17 71.53$\pm$.35 64.56$\pm$.27 76.86$\pm$.12
IV-C Results on ImageNet-1k
TABLE IX: Top-1 classification accuracy (%) on ImageNet-LT and Places-LT. * means results reported in the original paper.
Dataset Methods closed-set setting open-set setting
Many-shot Medium-shot Few-shot Overall Many-shot Medium-shot Few-shot F-measure
ImageNet-LT OLTR 43.2$\pm$0.1* 35.1$\pm$0.2* 18.5$\pm$0.1* 35.6$\pm$0.1* 41.9$\pm$0.1* 33.9$\pm$0.1* 17.4$\pm$0.2* 44.6$\pm$0.2*
OLTR+AdaAugment 45.9$\pm$0.1 38.3$\pm$0.1 22.0$\pm$0.2 39.0$\pm$0.1 44.1$\pm$0.1 36.8$\pm$0.1 20.8$\pm$0.2 45.8$\pm$0.1
Places-LT OLTR 44.7$\pm$0.1* 37.0$\pm$0.2* 25.3$\pm$0.1* 35.9$\pm$0.1* 44.6$\pm$0.1* 36.8$\pm$0.1* 25.2$\pm$0.2* 46.4$\pm$0.1*
OLTR+AdaAugment 43.7$\pm$0.1 41.1$\pm$0.1 29.5$\pm$0.2 39.6$\pm$0.1 43.9$\pm$0.1 40.8$\pm$0.1 28.9$\pm$0.1 50.4$\pm$0.1
We evaluate AdaAugment on the more challenging large-scale IamgeNet-1k dataset [36]. Note that the method AdvMask is not compared on ImageNet-1k due to its huge time costs of the sparse adversarial attack module [16]. As shown in Table VI and Table VII with more advanced baseline models, on ImageNet-1k, all DA methods’ average accuracy significantly outperforms the baseline. Moreover, AdaAugment achieves the highest improvements in accuracy using ResNet-50. Notably, AdaAugment achieves 1.1% improvements compared to the baseline without any architecture modifications and noticeable additional training costs. Although MADAug obtains similar results to ours on ImageNet-1k, it involves a bi-level optimization scheme that introduces much higher training costs than AdaAugment. Moreover, in contrast to previous works, which employed fixed augmentation operations for the entire dataset, our method applies model- and sample-adaptive operations for each image. This significantly enhances the model’s generalization.
IV-D Transfer Learning
@@
Transfer learning plays a pivotal role in assessing the models’ transferability [77], a common practice in data augmentation [30, 16, 9]. To quantitatively assess the efficacy of various DA methods, we assess the transferred accuracy of models using different DA methods. Specifically, we pre-train ResNet-50 models on CIFAR-100 and Tiny-ImageNet utilizing different augmentations, followed by fine-tuning these models on the CIFAR-10 dataset. Consequently, superior DA techniques result in higher transferred test accuracy.
@@
In Table VIII, we present the transferred accuracy achieved by various DA methods. While the discrepancies in transferred accuracy may appear modest, it is noteworthy that AdaAugment consistently outperforms others. This consistent and effective performance underscores our method’s capability to enhance model transferability performance. Thus, AdaAugment indeed enhances the model generalization performance.
IV-E Results on Long-Tail Datasets
While most DA methods have not been tested on large-scale long-tail biased datasets, we strengthen the generality and effectiveness of our method by applying it to the ImageNet-LT and Places-LT datasets [37]. Specifically, we utilize the codebase provided by OLTR [37] and rigorously follow all the hyperparameters and training strategies.
As shown in Table IX, AdaAugment brings substantial improvements across all the many/medium/few-shot classes and the open classes in both long-tail datasets. Notably, AdaAugment achieves over 3% accuracy improvements in the closed-set setting, highlighting the effectiveness of our method in improving model robustness and generalization. Therefore, AdaAugment can also be effectively employed in more challenging long-tail datasets as well.
IV-F Results on Fine-grained Datasets
TABLE X: Test accuracy (%) of AdaAugment on fine-grained datasets with ResNet-50. The performance is average±$\pm$±std.
Dataset baseline AdaAugment
Oxford Flowers [39] 89.47$\pm$0.08 97.17$\pm$0.14
Oxford-IIIT Pets [40] 89.73$\pm$0.18 91.95$\pm$0.24
FGVC-Aircraft [41] 77.25$\pm$0.09 90.92$\pm$0.05
Stanford Cars [38] 82.13$\pm$0.03 84.76$\pm$0.20
To comprehensively assess the effectiveness of our proposed methods, we apply AdaAugment to various fine-grained datasets, including Oxford Flowers [39], Oxford-IIIT Pets [40], FGVC-Aircraft [41], and Stanford Cars [38]. Specifically, for all the fine-grained datasets, we employ the ResNet-50 model [73] pre-trained on ImageNet, followed by fine-tuning these models using our proposed methods. To ensure fairness, experiments on the same dataset utilize the same experimental settings.
Table X summarizes the test accuracy of our methods compared to the baseline. It can be observed that AdaAugment brings notable accuracy improvements across all fine-grained datasets. Specifically, AdaAugment improves the model performance on Oxford Flowers by 7.70%, Oxford-IIIT Pets by 2.22%, FGVC-Aircraft by 13.67%, and Stanford Cars by 2.63%, respectively. Hence, our proposed methods can also be utilized to substantially enhance the model performance on fine-grained datasets.
IV-G Analytical Results
@@
Figure 4: Visualization of effectiveness using t-SNE algorithm with CIFAR-10 dataset. The embedding model is ResNet-50. DI: Dunn index.
Visualization of effectiveness. Since the primary objective of DA is to enhance the generalization performance of models, in this section, we conduct a comparative analysis of effectiveness visualization between models utilizing AdaAugment and those without. Specifically, we leverage the t-SNE algorithm [78] to visualize the feature embedding of the CIFAR-10 test set, utilizing feature maps derived from different trained models. Theoretically, models that showcase more robust generalization capabilities tend to excel in feature extraction ability.
Figure 4 presents the t-SNE visualization results for the baseline model and the model trained with AdaAugment. Compared to Figure 4LABEL:sub@fig2-1, it can be seen that Figure 4LABEL:sub@fig2-2 exhibits an optimized geometric structure characterized by enhanced inter-cluster separation and intra-cluster compactness. Furthermore, to furnish a quantitative evaluation of these embedding results, we employ the Dunn Index (DI) [79] as an evaluative metric. The DI is mathematically defined as follows:
$DI=\frac{\min_{1\leq i\neq j\leq m}\delta\left(C_{i},C_{j}\right)}{\max_{1\leq j\leq m}\Delta_{j}},$ (5)
where separation $\delta\left(C_{i},C_{j}\right)$ denotes the inter-cluster distance metric between clusters $C_{i}$ and $C_{j}$, and compactness $\Delta_{j}$ calculates the mean distance between all pairs within each cluster. A higher DI value signifies superior clustering performance. The DI values for the baseline in Figure 4LABEL:sub@fig2-1 and AdaAugment in Figure 4LABEL:sub@fig2-2 are $5.03\times 10^{-5}$ and $1.05\times 10^{-4}$, respectively. Consequently, AdaAugment achieves a DI value that is 108.7% higher than that of the baseline, demonstrating its superior efficacy in enhancing model performance. Through both quantitative and qualitative analyses, we validate the effectiveness of AdaAugment in improving model performance.
TABLE XI: The additional architecture parameters and training time of the auxiliary policy network on CIFAR-10. The device used is 2 NVIDIA RTX2080TI GPUs and an Intel(R) Xeon(R) CPU E5-2678 @ 2.50GHz.
@@
Complexity analysis.  Most existing DA methods typically rely on predefined random magnitude values during training. While this mechanism incurs minimal additional computational overhead in online training, it may not be optimally aligned with the dynamic evolution of deep models, potentially introducing undesirable side effects. To address this, AdaAugment incorporates an auxiliary policy network into online training to dynamically determine the magnitudes of DA operations. Although this adjustment inevitably introduces slightly additional parameters and computational complexity, it is a deliberate trade-off to achieve more effective DA.
@@
In this section, following the methodology employed in prior studies [21, 46, 24], we thoroughly analyze the parameter and time complexity of AdaAugment to assess its efficiency. Since the policy network’s parameter complexity is related to the feature space of the target network, we report the parameter complexity on ResNet-18/50 and WRN-28-10 in Table XI. Notably, it can be observed that the policy network’s parameter complexity results in a mere 1.3% increase for ResNet-18 (with 11.7M parameters), a 2.4% increase for ResNet-50 (with 25.5M parameters), and a 0.52% increase for WRN-28-10 (with 36.5M parameters). Furthermore, the overall training costs experience a slight increase for these networks. Despite the marginal increases in training costs, AdaAugment achieves significant accuracy improvements, with a notable enhancement of 1.47%, 1.68%, and 2.04% for ResNet-18, ResNet-50, and WRN-28-10, respectively. This improvement underscores AdaAugment’s remarkable effectiveness-efficiency trade-off, demonstrating its capacity to enhance model performance while incurring minimal training costs.
@@
Figure 5: Convergence analysis of training process with CIFAR-10 using ResNet-50.
Figure 6: Evolving magnitudes throughout the training process with CIFAR-100 using ResNet-50.
Convergence analysis.  In this section, we conduct a convergence analysis of AdaAugment throughout the entire training process, comparing it against two of the most representative DA methods, AutoAugment and TrivialAugment. Specifically, to illustrate the dynamic evolution of test errors during training, we train ResNet-18 models on CIFAR-10 using a multi-step learning rate decay schedule. The learning rate is initialized as 0.1 and multiplied by 0.2 at epochs 60, 120, 160, 220, and 280. From Figure 5, it can be observed that AdaAugment significantly enhances model performance, particularly after the second learning rate drop. Moreover, even after the initial learning rate drop, AdaAugment consistently maintains lower error rates than other methods. These empirical findings not only underscore the superior efficacy of AdaAugment in terms of model performance enhancement but also emphasize its capacity to expedite the convergence of the model during training. Therefore, AdaAugment contributes to decreasing the training difficulty of deep networks.
Adaptively evolving magnitude. To present the evolving progress of the adaptive magnitudes, we train ResNet-50 with CIFAR-100. The learning rate is initialized as 0.1 and multiplied by 0.2 at epochs 60, 120, 160, 220, and 280, along with a 5-epoch scheduler warmup. In Figure 6, it can be seen that the magnitude generally increases through model training, providing progressively more diverse augmented samples to mitigate the risk of overfitting. Moreover, this dynamic alignment aligns with the model’s training status. For instance, at different training stages, such as the warmup phase at the beginning of training or during learning rate changes, the magnitudes adaptively modify values, aligning with the model’s evolving status.
(a) Training Accuracy.
(b) Test Accuracy.
Figure 7: Effectiveness of AdaAugment in mitigating overfitting risks with reduced CIFAR-10 using ResNet-50.
TABLE XII: Impact of the augmentation space with CIFAR-10 using ResNet-18.
# of transformations 2 4 6 8 10 12 14
Accuracy (%) 95.99$\pm$0.02 96.18$\pm$0.04 96.38$\pm$0.05 96.54$\pm$0.10 96.61$\pm$0.11 96.67$\pm$0.09 96.75$\pm$0.06
Effectiveness of mitigating overfitting.  Although modern deep models can fit the distribution of training datasets through training, this capability often faces overfitting risks, which degrades models’ generalization performance, particularly in scenarios with limited training data. Data augmentation is primarily designed to address this issue by increasing the diversity of training data. To evaluate the effectiveness of AdaAugment in mitigating overfitting risks, we simulate a high-overfitting-risk scenario by training a large ResNet-50 model on the relatively simple CIFAR-10 dataset. To further amplify the overfitting risk, we randomly reduce the sizes of the training set by 50% while keeping the test set unchanged.
As can be seen in Figure 7, without AdaAugment, models achieve very high training accuracy, nearing 100%, but the test accuracy rises slowly and exhibits significant fluctuations, indicating poor generalization and significant overfitting. With AdaAugment, models achieve slower training accuracy growth and consistently lower training accuracy. Despite this, the test accuracy is significantly higher and more stable, demonstrating AdaAugment’s ability to mitigate overfitting and enhance generalization performance.
TABLE XIII: Impact of adaptive magnitude $m$. Comparative analysis of AdaAugment compared to different settings of $m$ with CIFAR-100 using ResNet-18/50.
Model $m=0.5$ random $m$ linear $m$ sine $m$ AdaAugment
R-18 78.62$\pm$0.32 77.08$\pm$0.30 78.38$\pm$0.27 78.58$\pm$0.32 79.84$\pm$0.27
R-50 80.23$\pm$0.29 80.61$\pm$0.19 80.28$\pm$0.31 80.65$\pm$0.15 81.46$\pm$0.12
TABLE XIV: Effect of different RL modules with ResNet-18 on CIFAR-10/100.
Dataset DDPG [62] SAC [61] Ours
CIFAR-10 96.73$\pm$0.06 96.82$\pm$0.04 96.75$\pm$0.06
CIFAR-100 79.27$\pm$0.19 79.80$\pm$0.29 79.84$\pm$0.27
TABLE XV: Effect of $\gamma$ with CIFAR-100 using ResNet-18/50.
$\gamma$ 0.2 0.4 0.6 0.8 0.99
R-18 78.94$\pm$0.33 78.95$\pm$0.20 78.99$\pm$0.29 79.24$\pm$0.25 79.84$\pm$0.27
R-50 80.61$\pm$0.15 80.44$\pm$0.20 80.68$\pm$0.17 80.90$\pm$0.19 81.46$\pm$0.12
TABLE XVI: Effect of $\lambda$ with CIFAR-100 using ResNet-18/50.
$\lambda$ 0 1 Ours
ResNet-18 78.53$\pm$0.36 78.70$\pm$0.30 79.84$\pm$0.27
ResNet-50 80.21$\pm$0.19 80.74$\pm$0.29 81.46$\pm$0.12
IV-H Ablation Study
Effect of the augmentation space. Although the augmentation space used in AdaAugment is not larger than other methods [21, 23, 22, 50, 46], we further conduct analytical experiments to evaluate the impact of base augmentations. Specifically, we analyze the performance by changing the number of augmentation transformations in our augmentation space. We present the results in Table XII. Notably, the number of augmentation types used in Table XII is consistently lower than other baseline approaches. It can be observed that while the accuracy decreases as fewer augmentations are utilized, it drops very slowly. When $N=2$, only two types of augmentations will be employed, which limits the diversity of augmented data. Even under this case, the accuracy drops less than 0.8%. This highlights the effectiveness of AdaAugment in adaptively adjusting the augmentation magnitudes, effectively enhancing the model generalization.
Effect of the RL module. By utilizing the RL module to adaptively adjust the augmentation magnitude of DA operations, AdaAugment eliminates the need for manual parameter design and adjustment. In this section, we conduct experiments to assess the effect of the RL module. By using the same augmentation space, we employ different experimental settings on $m$: fixed and random values of $m$, as well as increasing $m$ using linear and sine functions. The results are summarized in Table XIII. It can be seen that the RL module brings substantial improvements in model performance. By estimating the real-time model training status through the RL module, AdaAugment can optimize the augmentation magnitudes during training. Thus, we demonstrate that the superior efficacy is brought by the adaptive adjustment of augmentation magnitude rather than the base augmentation space.
Additionally, in AdaAugment, we utilize lightweight A2C architectures. Here, we also employ two other more recent RL architectures, including Soft Actor-Critic (SAC) [61] and Deep Deterministic Policy Gradient (DDPG) [62]. As can be seen in Table XIV, the performance differences among different RL modules are relatively marginal.
Effect of the discount factor $\gamma$. Following previous works [70, 71, 72], we set $\gamma$ as a constant 0.99. Here, we investigate the effect of different $\gamma$ values. The results are presented in Table XV. It can be seen that AdaAugment performs best under our setting.
Effect of parameter $\lambda$ in Eq. (2). We conduct analytical experiments to evaluate the effect of parameter $\lambda$ in Eq. (2). Specifically, instead of dynamically adjusting $\lambda$, $\lambda$ is fixed as 0 and 1. The results are shown in Table XVI, validating the stability of parameter $\lambda$.
V Discussion and Future Work
In this work, we propose AdaAugment, a tunning-free adaptive data augmentation method. Instead of relying on any hand-crafted criteria, AdaAugment incorporates an RL module to adaptively adjust the augmentation magnitudes for individual training samples. This adaptive adjustment aligns with the model training progress. Experiment results demonstrate that AdaAugment achieves enhanced model performance across various benchmark datasets without introducing noticeable training overhead. Besides strengths, the limitations of our method should also be discussed.
While the proposed method estimates the risks of underfitting and overfitting by deriving losses from fully augmented and nonaugmented data, this additional loss calculation requires two extra forward passes compared to vanilla training. To enhance the overall efficiency, our policy network is designed to be consistent across various target models. Thus, the increase in GPU hours is relatively modest and not directly correlated with the complexity of the task model. Future work should consider leveraging other surrogate indicators to estimate underfitting and overfitting risks.
In our augmentation space, we leverage image transformations whose applied strengths can be adjusted with a real-valued parameter, i.e., $m\in[0,1]$. However, most existing DA methods, such as Cutout [14], AdvMask [16], and AutoAugment [21], do not support explicit adjustments of augmentation magnitudes. Therefore, future work should consider exploring an adjustable version of existing advanced DA methods, facilitating explicit adjustment of augmentation strength with continuous real values. Consequently, our augmentation space may be further expanded. This expansion holds promise for further performance enhancements.
VI Conclusion